# Autonomous Imagination: Closed-Loop Decomposition of Visual-to-Textual Conversion in Visual Reasoning for Multimodal Large Language Models

### v3

@@ Abstract
Under pure textual modality, Large Language Models (LLMs) have demonstrated remarkable success in complex reasoning tasks by decomposing them into simpler sub-problems. However, Multimodal Large Language Models (MLLMs) still struggle with some seemingly straightforward visual tasks, such as counting and solving jigsaw puzzles. We argue that these tasks challenge the ability of visual-to-textual conversion, where MLLMs convert visual information perceived from the input scene, to textual information for further reasoning and generating the answer. If the complexity of the visual input is beyond the perceptual capability of the MLLMs, without decomposing this conversion process, simply scaling inference-time reasoning cannot solve the task because it repeatedly encounters the same perceptual bottleneck. We propose an approach, autonomous imagination, to enable MLLMs to iteratively modify visual inputs (e.g. isolating objects, rearranging puzzle pieces) into intermediate visual states, decomposing visual-to-textual conversion into closed-loop visual modification steps. We show that, without any retraining, MLLMs can now solve tasks initially beyond their perceptual capability, highlighting that closed-loop visual modification can be an effective way of decomposing the visual reasoning task into solvable substeps. Project page: https://future-item.github.io/autoimagine-site/.
@@ 2.2 Visual Reasoning in MLLMs
In addition to reasoning within text modalities, substantial efforts are made in reasoning in visual tasks. Existing MLLMs often face limitations in direct visual perception and are prone to generating answers with hallucinations (Zhang et al., 2023d). Initial approaches used attention mechanisms to improve question-answering abilities (Zhou et al., 2021), and subsequent solutions include strengthening reasoning in the visual modality by training (Lin et al., 2023; Wang et al., 2023b) or employing auxiliary knowledge (Mitra et al., 2023). In addition, researchers have proposed effective visual prompting methods to refine the focus of MLLMs (Jiang et al., 2024; Lin et al., 2024; Hong et al., 2023; Zhang et al., 2023b; c; Mitra et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Yu et al., 2024). See Wu et al. (2024a) for the recent comprehensive survey. Attempts have been made to utilize visual prompts in the Chain-of-Thought (CoT) paradigm (Wei et al., 2022). Based on pioneer studies (Zhang et al., 2023d; Zheng et al., 2023; Zhang et al., 2023a; Peng et al., 2023), recent approaches build CoT by enabling MLLMs to add visual prompts autonomously either through direct model training (Lai et al., 2023; Wu & Xie, 2023; Chen et al., 2023; Qi et al., 2024; Shao et al., 2024) or by utilizing external visual processing models (Zhou et al., 2024b). By using visual prompts (e.g., bounding boxes) to highlight key information in images, the visual-to-textual conversion step becomes easier, leading to a performance boost. However, even though visual prompts simplify the perception of input images, this paradigm still restricts to do visual-to-textual conversion only in one step, which is after all visual prompts are added in the input. As our experiments demonstrate, even when visual prompting is applied, visual-to-textual conversion remains a bottleneck in common visual reasoning tasks that are intuitive to humans, highlighting the necessity of effective task decomposition.
@@ 3.4 Focus Operator
To enable the focus operator in 3D Gaussian-based representations, we introduce a method to selectively segment an object from an unstructured 3D Gaussian scene, which is illustrated in Alg. 1. Given the MLLM’s initial selection block, we first generate a virtual camera trajectory that orbits around the targeted object, creating a video sequence. This sequence is fed into the SAM2 model (Ravi et al., 2024), which applies the selection cursor condition on all rendered frames. For each frame, rays are cast from pixels within the segmentation mask, and path tracing records the radiance contributions of intersecting Gaussians. This process is expressed as volumetric integration illustrated in Line 4-11. Gaussians who receive a contribution higher than a threshold will receive one vote.
@@ 12            $T\leftarrow T\cdot(1-\alpha_{i})$
21       if votes[$g_{i}$] $\geq$ 0.1×0.1$\times$0.1 × max_vote then
@@ 4.1 Benchmark and Evaluation Protocol
Multi-Object Hallucination: Current MLLMs can suffer from hallucination by perceiving or generating non-existing objects in the input scene. The complexity of the input scene is beyond the limitation in perceptual capability of MLLMs, which is a plausible reason for hallucination. Our approach can serve as a promising way to address this challenge, especially for the scenes where multiple objects exist. By the closed-loop iterative process of autonomous imagination, MLLMs can deal with objects one by one, as in our counting benchmark, instead of handling them all together, effectively reducing the possibility of hallucination. We conducted additional experiments on the recently proposed multi-object hallucination benchmark ROPE (Chen et al., 2024) to validate this argument. We focus on the most challenging subset in the benchmark where the state-of-the-art model on the official leaderboard, GPT-4o, exhibits the poorest performance, which is answering about single object given multiobject image, heterogeneous object types, on unseen data split. We refer to the detailed benchmark setting from the original paper.
@@ 4.6 Multi-Object Hallucination Benchmark
We evaluated our method using the recently proposed ROPE benchmark for Multi-Object Hallucination (Chen et al., 2024), focusing on the most challenging subset where GPT-4o exhibits the poorest performance, which is answering about a single object given multi-object image, heterogeneous object types, on unseen data split. Our method uses GPT-4o as the base model that equips the imagination space. Please refer to the original paper for more details. Given that this benchmark emphasizes the identification of abstract objects that are not suitable for segmentation, we turn the focus operation into drawing a large rectangular region by MLLMs natively, through selecting top-left and bottom-right corners. The benchmark indicates that MLLMs are prone to hallucinations when confronted with multiple objects that introduce additional visual distractions. We demonstrate that using our method, MLLMs can autonomously focus on the important region despite these distractions, leading to more accurate responses due to the elimination of irrelevant visual information.
[…]
@@
4.7 Discussions
Time cost. The time cost of our approach depends mainly on three factors: the running time of a single call of imagination operators, a single call of MLLM reasoning, and the number of MLLM calls. The calls of imagination operators are quite efficient. For example, each call of the SAM model takes 220ms on NVIDIA RTX 4090 on average, which is very efficient compared with MLLM reasoning. Considering the number of calls for MLLM reasoning, the average numbers of MLLM API call for our approach are: counting=117 (for all objects), jigsaw puzzle=20 (for one piece), object placement=15 (for one object), multi-object hallucination=26 (for one object). As shown by the comparisons on GPT-4o Sampling baseline in Tab. 1, our approach is more efficient than previous sampling-based imagination strategies (Kapelyukh et al., 2024; Ding et al., 2024), the closest previous studies to ours, by enabling to solve broader tasks with vast solution spaces.
@@ 5 Conclusion
Limitations and future work. Our work is a proof of concept: while closed-loop task decomposition overcomes the visual-to-textual conversion bottleneck in our demonstrated tasks, resource constraints prevented us from training a model with native closed-loop reasoning and visual modification capabilities. This also prevents us from validating our approach under large, general-purpose visual benchmarks. We hope that our work inspires future efforts to close this critical gap.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (U23A20311,62206245).
Appendix A More Experiments
@@
Appendix B Additional Qualitative Results
@@
Appendix C System Prompts
We provide system prompts for the reasoning instructions given to the MLLMs in our tasks, as shown in Fig. 15, as well as the prompts used as reminders for autonomous imagination operations, as illustrated in Fig. 16.

### v4

@@ Abstract
Under pure textual modality, Large Language Models (LLMs) have demonstrated remarkable success in complex reasoning tasks by decomposing them into simpler sub-problems. However, Multimodal Large Language Models (MLLMs) still struggle with some seemingly straightforward visual tasks, such as counting and solving jigsaw puzzles. We argue that these tasks challenge the ability of visual-to-textual conversion, where MLLMs convert visual information perceived from the input scene, to textual information for further reasoning and generating the answer. If the complexity of the visual input is beyond the perceptual capability of the MLLMs, without decomposing this conversion process, simply scaling inference-time reasoning cannot solve the task because it repeatedly encounters the same perceptual bottleneck. We propose an approach, autonomous imagination, to enable MLLMs to iteratively modify visual inputs (e.g. isolating objects, rearranging puzzle pieces) into intermediate visual states, decomposing visual-to-textual conversion into closed-loop visual modification steps. We show that, without any retraining, MLLMs can now solve tasks initially beyond their perceptual capability, highlighting that closed-loop visual modification can be an effective way of decomposing the visual reasoning task into solvable substeps. Our code and data are released at https://future-item.github.io/autoimagine-site/.
@@ 2.2 Visual Reasoning in MLLMs
In addition to reasoning within text modalities, substantial efforts are made in reasoning in visual tasks. Existing MLLMs often face limitations in direct visual perception and are prone to generating answers with hallucinations (Zhang et al., 2023d). Initial approaches used attention mechanisms to improve question-answering abilities (Zhou et al., 2021), and subsequent solutions include strengthening reasoning in the visual modality by training (Lin et al., 2023; Wang et al., 2023b), employing auxiliary knowledge (Mitra et al., 2023), and utilizing external visual perception modules Surís et al. (2023). In addition, researchers have proposed effective visual prompting methods to refine the focus of MLLMs (Jiang et al., 2024; Lin et al., 2024; Hong et al., 2023; Zhang et al., 2023b; c; Mitra et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Yu et al., 2024). See Wu et al. (2024a) for the recent comprehensive survey. Attempts have been made to utilize visual prompts in the Chain-of-Thought (CoT) paradigm (Wei et al., 2022). Based on pioneer studies (Zhang et al., 2023d; Zheng et al., 2023; Zhang et al., 2023a; Peng et al., 2023), recent approaches build CoT by enabling MLLMs to add visual prompts autonomously either through direct model training (Lai et al., 2023; Wu & Xie, 2023; Chen et al., 2023; Qi et al., 2024; Shao et al., 2024) or by utilizing external visual processing models (Zhou et al., 2024b). By using visual prompts (e.g., bounding boxes) to highlight key information in images, the visual-to-textual conversion step becomes easier, leading to a performance boost. However, even though visual prompts simplify the perception of input images, this paradigm still restricts to do visual-to-textual conversion only in one step, which is after all visual prompts are added in the input. As our experiments demonstrate, even when visual prompting is applied, visual-to-textual conversion remains a bottleneck in common visual reasoning tasks that are intuitive to humans, highlighting the necessity of effective task decomposition.
@@
Recently, advanced image editing models that accept input of natural languages are proposed (Huang et al., 2024; Wang et al., 2024). Notably, concurrent work has enables image editing models with strong reasoning abilities (Fang et al., 2025). These models have the potential to be utilized as the imagination tools in our approach, while are more capable to deal with visual generation and editing tasks by themselves.
@@ 3.4 Focus Operator
To enable the focus operator in 3D Gaussian-based representations, we introduce a method to selectively segment an object from an unstructured 3D Gaussian scene [^1] We refer the detailed introduction of the 3D conditional segmentation algorithm to our open-released code., which is illustrated in Alg. 1. Given the MLLM’s initial selection block, we first generate a virtual camera trajectory that orbits around the targeted object, creating a video sequence. This sequence is fed into the SAM2 model (Ravi et al., 2024), which applies the selection cursor condition on all rendered frames. For each frame, rays are cast from pixels within the segmentation mask, and path tracing records the radiance contributions of intersecting Gaussians. This process is expressed as volumetric integration illustrated in Line 4-11. Gaussians who receive a contribution higher than a threshold will receive one vote.
@@ 12    $T\leftarrow T\cdot(1-\alpha_{i})$
21 if votes[$g_{i}$] $\geq$ $0.1\times$ max_vote then
@@ 3.7 Reasoning Process
3.8 System Prompts
Our approach is training-free and utilizes the native reasoning ability of the MLLMs to autonomously utilize the visual imagination operators provided. We design a set of general system prompts across all tasks. The reasoning instructions given to the MLLMs in our tasks are shown in Fig. 15. The prompts used as reminders for autonomous imagination operations are illustrated in Fig. 16.
@@ 4.1 Benchmark and Evaluation Protocol
Multi-object hallucination: Current MLLMs can suffer from hallucination by perceiving or generating non-existing objects in the input scene. The complexity of the input scene is beyond the limitation in perceptual capability of MLLMs, which is a plausible reason for hallucination. Our approach can serve as a promising way to address this challenge, especially for the scenes where multiple objects exist. By the closed-loop iterative process of autonomous imagination, MLLMs can deal with objects one by one, as in our counting benchmark, instead of handling them all together, effectively reducing the possibility of hallucination. We conducted additional experiments on the recently proposed multi-object hallucination benchmark ROPE (Chen et al., 2024a) to validate this argument. We focus on the most challenging subset in the benchmark where the state-of-the-art model on the official leaderboard, GPT-4o, exhibits the poorest performance, which is answering about single object given multiobject image, heterogeneous object types, on unseen data split. We refer to the detailed benchmark setting from the original paper.
@@ 4.6 Multi-Object Hallucination Benchmark
We evaluated our method using the recently proposed ROPE benchmark for Multi-Object Hallucination (Chen et al., 2024a), focusing on the most challenging subset where GPT-4o exhibits the poorest performance, which is answering about a single object given multi-object image, heterogeneous object types, on unseen data split. Our method uses GPT-4o as the base model that equips the imagination space. Please refer to the original paper for more details. Given that this benchmark emphasizes the identification of abstract objects that are not suitable for segmentation, we turn the focus operation into drawing a large rectangular region by MLLMs natively, through selecting top-left and bottom-right corners. The benchmark indicates that MLLMs are prone to hallucinations when confronted with multiple objects that introduce additional visual distractions. We demonstrate that using our method, MLLMs can autonomously focus on the important region despite these distractions, leading to more accurate responses due to the elimination of irrelevant visual information.
Table 2: Results under multi-object hallucination benchmark. Except for o1, Molmo, and our method, the results of other baselines are cited from the official leaderboard in https://multi-object-hallucination.github.io/. For each base model family, we take the one with the highest performance. Mechanistically grounded MLLMs (marked with *) take visual prompts by dedicated pointer tokens. Please refer to the original paper for more details.
@@
4.7 Discussions
Table 3: Results under more existing benchmarks.
Models CLEVR (Counting) Where2Place (Object Placement)
Success Rate Mean Error Variance Accuracy
GPT-4o 57.3% 0.52 0.45 29.1%
GPT-4o + ours 74.7% 0.42 0.74 37.0%
Table 4: Results for smaller MLLMs under the object placement task.
Qwen2.5-VL-7B Qwen2.5-VL-32B InternVL2_5-8B InternVL2_5-26B
Base Model Ours Base Model Ours Base Model Ours Base Model Ours
Locating 7.2% 19.3% 2.9% 16.8% 1.1% 11.6% 1.5% 16.7%
Placement 1.1% 2.2% 0.85% 4.2% 0% 3.6% 0.7% 5.1%
Table 5: Comparison between RobotPoint (fine-tuned to solve Where2Place tasks) and GPT-4o enhanced with our approach.
Models Counting Simple Jigsaw Puzzle Object Placement (Ours) Object Placement (Where2Place)
Success Rate Mean Error Variance 4-Piece Missing 6-Piece Missing Locating Placing Accuracy
RobotPoint 25.0% 2.04 3.92 6.8% 6.1% 1.0% 0.03% 46.8%
GPT-4o 39.8% 0.73 0.62 29.5% 9.1% 10.9% 3.6% 29.1%
GPT-4o + ours 85.3% 0.19 0.22 68.2% 51.5% 69.4% 37.3% 37.0%
Table 6: Comparison between OpenAI’s o1 and o3 models under our benchmarks.
Models Counting Simple Jigsaw Puzzle Object Placement
Success Rate Mean Error Variance 4-Piece Missing 6-Piece Missing Locating Placing
o1 50.8% 0.74 0.90 31.8% 27.3% 17.3% 8.5%
o3 w/o tool 50.0% 0.71 0.78 61.4% 66.7% 9.5% 4.4%
o3 76.2% 0.36 0.69 78.6% 57.1% 22.9% 6.2%
Results on more existing benchmarks. To verify the robustness of our approach, we further conduct experiments under two more existing benchmarks. One is CLEVR (Johnson et al., 2017) for the counting task. In each image from CLEVR, a scene containing various number of geometric objects is presented. We randomly sample from the official test set to ensure balance over the numbers of objects in the images, which includes 178 images. Another is Where2Place (Yuan et al., 2024), which includes 100 real-world scenes from homes and offices in the wild for the object placement task. The results are shown in Tab. 3, verifying that our approach consistently improves the reasoning ability of the base MLLM.
Applying on smaller-sized MLLMs. Our approach requires the ability of the base MLLMs to autonomously utilize the visual imagination operators provided. This is natively satisfied by larger state-of-the-art MLLMs, such as GPT-4o, while maybe challenging for smaller-sized MLLMs without specific training. We verify this in the counting and jigsaw puzzle solving tasks, where the smaller-sized MLLMs usually have the difficulty in realizing “when to stop”: Common issues include incorrectly deciding to stop counting when there still exist remaining objects, or being hard to judge whether the jigsaw piece has been put at the right position. This shows the essentialness of studying how the autonomous imagination ability can be trained, which is left as future work. On the other hand, we discover that our training-free strategy can indeed be utilized to enhance smaller-sized MLLMs (Qwen2.5 (Bai et al., 2025) and InternVL2.5 (Chen et al., 2024b)) under the object placement task, which is shown in Tab. 4, verifying the potential of our approach for smaller-sized MLLMs.
Comparing with in-domain training. It is also meaningful to compare our approach with in-domain trained models. In Tab. 1, we have shown that Molmo achieves strong performance in the counting task, with the specifically trained ability to utilize visual cues in counting, while performing not ideally under other tasks. We further test the RobotPoint model (Yuan et al., 2024), trained specifically for predicting image keypoint affordances given language instructions, which is especially suitable to handle the Where2Place benchmark. We compare its performance under Where2Place, as well as our counting, jigsaw puzzle solving, and object placement tasks, as shown in Tab. 5. Similar to Molmo, RobotPoint shows strong performance under Where2Place that it is trained to handle, while fails to generalize its ability in other tasks. This suggests that in-domain training is more suitable to address specific tasks, while improving test-time reasoning can be an easier way to improve general reasoning ability across different tasks. Moreover, as discussed in the analysis of smaller-sized MLLMs, conducting training on the general reasoning ability instead of domain-specific task solving is also promising.
Concurrent work on multimodal CoT reasoning. Concurrent to our work, there is a surge of research for multimodal CoT reasoning. A representative is OpenAI’s o3 model (OpenAI, 2025a), which is natively trained multimodal reasoning ability for “thinking with images” (OpenAI, 2025b), which utilizes visual processing tools to modify images (e.g. crop, zoom, rotate), hence realize more thorough analysis of the inputs during CoT reasoning. This paradigm is also studied by other works via visual test-time scaling (Luo et al., 2025; Wu et al., 2025), supervised fine-tuning (Li et al., 2025; Fu et al., 2025), and reinforcement learning (Duan et al., 2025; Xu et al., 2025). Tab. 6 illustrates the comparison among OpenAI’s o1 and o3 models in our proposed benchmarks, with an additional baseline “o3 w/o tool” in which we explicitly prompt o3 to disable the tool use ability. The results are consistent with our work: Multimodal CoT usually improves over purely textual reasoning. Compared with these concurrent works, our work uniquely proposes the notion of “visual-to-textual” conversion and explicitly differentiates between visual prompting and visual imagination. This can lead to better understanding of a fundamental research question: In what kinds of tasks, the ability to think with images is necessary? Furthermore, our plug-and-play autonomous imagination approach can be of independent interest to inspire more advanced multimodal CoT reasoning approaches.
Time cost. The time cost of our approach depends mainly on three factors: the running time of a single call of imagination operators, a single call of MLLM reasoning, and the number of MLLM calls. The calls of imagination operators are quite efficient. For example, each call of the SAM model takes 220ms on NVIDIA RTX 4090 on average, which is very efficient compared with MLLM reasoning. Considering the number of calls for MLLM reasoning, the average numbers of MLLM API call for our approach are: counting=117 (for all objects), jigsaw puzzle=20 (for one piece), object placement=15 (for one object), multi-object hallucination=26 (for one object). As shown by the comparisons on GPT-4o Sampling baseline in Tab. 1, our approach is more efficient than the closest previous studies to ours: sampling-based imagination strategies (Kapelyukh et al., 2024; Ding et al., 2024), by enabling to solve broader tasks with vast solution spaces.
@@ 5 Conclusion
Limitations and future work. Our work is a proof of concept: While closed-loop task decomposition overcomes the visual-to-textual conversion bottleneck in our demonstrated tasks, resource constraints prevented us from training a model with native closed-loop reasoning and visual modification capabilities. This also prevents us from validating our approach under large, general-purpose visual benchmarks. We hope that our work inspires future efforts to close this critical gap. Furthermore, even though we test the robustness of our approach in counting and jigsaw puzzle sovling through different levels of difficulty (in terms of numbers of objects/pieces), for object placement and multi-object hallucination tasks, properly measuring the task difficulty is itself a challenging problem. These tasks involve real-world scenes with complicated spatial and semantical relationships among the objects, which may not be easily captured by a single measurement of hardness. We treat this as an important future research problem. Finally, our current approach limits to utilize chain reasoning with fixed plans, which is not optimal in terms of time cost. Future work can improve the efficiency of reasoning with more advanced searching and planning strategies.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (U23A20311,62206245). We would like to thank the reviewers for their insightful and constructive suggestions.
Appendix A Potential Broader Impact
Our work is largely foundational and aims to improve the reasoning capabilities and reliability of multimodal language models. Reducing hallucinations in complex visual scenes could enhance safety and robustness in real-world applications. On the other hand, our approach could increase the computational cost of MLLMs, leading to energy inefficiency and environmental impact. Reproducing our work also requires access to existing MLLMs and visual tools, leading to additional complexity and costs. We hope that our efforts on open-release all our code and data could reduce this burden.
Appendix B More Experiments
@@
Appendix C Additional Qualitative Results